OpenStudy (anonymous):

Please help: Eliminate the parameter. x = 3 cos t, y = 3 sin t

OpenStudy (anonymous):

x^2=9(cosx)^2 y^2=9(sinx)^2 x^2+y^2=9(cos)^2+9(sinx)^2 x^2+y^2=9

OpenStudy (anonymous):

(cosx)^2+(sinx)^2=1

OpenStudy (anonymous):

you have a circle

OpenStudy (anonymous):

Thanks kitty, so helpful. i might need help on a different type of question soon...